The viewing window of a calculator has a perimeter of 20 centimeters and is 2 centimeters longer than it is wide. Find the dimensions of the window
Answer by rimjhim.tomar(21)

let the width be x cm.
then by the given information we have length= x+2 cm
hence,
the perimeter of the rectangular window=20 cm
2(length+width) = 20
2(x+x+2) =20
2(2x+2) =20
dividing by 2 on both sides
2x+2 =20
2x =20
x =10
hence the width is x=10 cm
and length=x+2= 10+2 cm =12 cm